If the equation for a market demand curve is Qd = 15 – 0,2P and the equation for a market supply curve is Qs = ̶ 1 + 0,6P, the market equilibrium price and quantity are? 


1
Expert's answer
2022-03-15T12:07:14-0400

Given that

Qd= 15 - 0.2p and

Qs= -1+0.6p


E=Qd=Qs=0

"15-0.2p=-1+0.6p"

"16=0.8p"

"P^*=20"

"Q^*=15-0.2(20)=11"